# A fast neural model for Bayesian Mixed-Effects Regression

### v1

@@ 2.4 Learning
To calculate the loss for the global parameters, we use the forward Kullback-Leibler divergence between the unknown true posterior $p(\bm{\vartheta}|\mathbf{s})$ and its flow-based approximation $p_{\Pi}(\bm{\vartheta}|\mathbf{s}):=p_{n}(\mathbf{z}_{n}|\mathbf{s})$,
$\ell_{\Pi}(\bm{\vartheta},\mathbf{s})\propto-\mathbb{E}_{\bm{\vartheta},\mathbf{s}}\left[\log p_{\Pi}(\bm{\vartheta}|\mathbf{s})\right]=-\mathbb{E}_{\bm{\vartheta},\mathbf{s}}\left[\log p_{0}(T(\bm{\vartheta}|\mathbf{s}))+\log\begin{vmatrix}\det J_{T}(\bm{\vartheta}|\mathbf{s})\end{vmatrix}\right$],
where $T$ and thereby the approximation $p_{\Pi}(\bm{\vartheta}|\mathbf{s})$ depend on the posterior network. Since the summary $\mathbf{s}$ of data $\mathbf{D}$ is itself depending on the summary network, the end-to-end loss can be written as
$\ell_{\Pi,\Sigma}(\bm{\vartheta},\mathbf{D})\propto-\mathbb{E}_{\bm{\vartheta},\mathbf{D}}\left[\log p_{\Pi}(\bm{\vartheta}|f_{\Sigma}(\mathbf{D}))\right]=-\mathbb{E}_{\bm{\vartheta},\mathbf{D}}\left[\log p_{0}(T(\bm{\vartheta}|f_{\Sigma}(\mathbf{D}))+\log\begin{vmatrix}\det J_{T}(\bm{\vartheta}|f_{\Sigma}(\mathbf{D}))\end{vmatrix}\right$].
The objective for the local parameters is completely analogous. We sum the local losses over groups and add the result to the global loss, which follows a potential factorization of the joint posterior over both types of regression parameters (see Appendix D). The expectation is approximated by averaging over the batch. Model weights are updated using the Schedule-Free AdamW optimizer (Defazio et al., 2024). We train separate models for different numbers of fixed effects and random effects until convergence, which requires between $10^{5}$ and $10^{6}$ training sets in our case.
@@ 2.5 Post-Hoc Refinement
Learning $p_{\Pi}$ by minimizing the forward Kullback-Leibler Divergence naturally forces $p_{\Pi}$ to be positive wherever $p$ is positive, making $p_{\Pi}$ mass-covering (Jerfel et al., 2021). Thus, we can use importance sampling to improve posterior estimation (Tokdar & Kass, 2010; Dax et al., 2023). For each sample $\bm{\vartheta}_{k}\sim p_{\Pi}(\bm{\vartheta}|\mathbf{D})$ we assign an importance weight,
$w_{k}=\frac{p(\mathbf{D}|\bm{\vartheta}_{k})p(\bm{\vartheta}_{k})}{p_{\Pi}(\bm{\vartheta}_{k}|\mathbf{D})},$
@@
Calibration with Conformal Prediction
Uncertainty quantification is a hallmark of Bayesian inference, making the fidelity of the approximate posterior’s credible intervals a critical concern. Posterior samples can be used to calculate empirical quantiles and thus also intervals that contain $c\%$ of the posterior density. Due to the mass-covering property of $p_{\Pi}$, the learned posteriors tend to be too wide – i.e. the true parameter is inside the $c\%$ credible interval in more than $c\%$ of the cases. This is a commonly known issue of normalizing flows (Chen et al., 2025; Dheur & Taieb, 2025). Conformal prediction (Vovk et al., 2022; Shafer & Vovk, 2008; Angelopoulos & Bates, 2022) is a general-purpose method that constructs distribution-free prediction sets $\hat{C}_{\alpha}$ such that $\mathbb{P}(\bm{\vartheta}\in\hat{C}_{\alpha})\geq 1-\alpha$. To construct $\hat{C}_{\alpha}$, we use a calibration set to calculate the difference between the true $\bm{\vartheta}$ and the closest border of the proposed $1-\alpha$ credible interval $C_{\alpha}$. The $1-\alpha$ quantile of these differences is then added to the proposed interval borders, widening them if the value is positive and narrowing them otherwise. Importantly, this does not require retraining but efficiently refines credible intervals post-hoc.
@@ 3 Results
We test our model against HMC on a toy dataset with highly constrained parameters and uncorrelated normal data (Section 3.1), in-distribution on held-out synthetic test sets of varying difficulty (Section 3.2), and out-of-distribution on semi-synthetic data where predictors $\mathbf{X}$ are taken from four real datasets (Section 3.3). This approach has the following considerable benefits over evaluation on purely real data: (1) The regression models are always correctly specified, (2) we know the ground truth parameters and can thus evaluate parameter recovery and coverage, (3) we can compare the results to in-distribution test data and gauge how well the model transfers to realistic predictors (Lueckmann et al., 2021; Ward et al., 2022). We use the following evaluation metrics: We quantify parameter recovery with Pearson’s correlation $r$ and $\mathrm{RMSE}$ between the true parameters and posterior means. We check the fidelity of credible intervals using coverage errors: A $\mathrm{CE}(\alpha)=0.05$ means that the model’s $1-\alpha$ credible interval $C_{\alpha}$ is on average 5% too wide. Coverage error are estimated using the difference between the relative frequency of the true parameter being inside $C_{\alpha}$ and the credible interval’s nominal probability mass ($1-\alpha$):
$\mathrm{CE(\alpha)}=\frac{1}{B}\sum_{b=1}^{B}\mathbbm{1}\left(\vartheta^{(b)}\in C_{\alpha}^{(b)}\right)-(1-\alpha).$
Finally, we plot posterior predictive distributions (Gelman et al., 2013), which visualize how much the posterior predictive samples $\mathbf{\tilde{y}_{t}}\sim p(\mathbf{y}|\bm{\hat{\vartheta}_{t}})$ match the actual $\mathbf{y}$. All metrics are calculated for both metabeta and HMC posterior samples. A brief comparison of required computation time is reported in Figure 1D.
@@ 3.1 Toy Example
To gauge if the pipeline works for both our model and HMC, we first test both on a toy example with $d=2$ and $q=1$. Regression parameters and observed data all fall inside small ranges, and the observed single predictor is sampled from a standard normal distribution. Result figures can be found in Appendix F. Both models reach almost perfect parameter recovery correlation for fixed effects ($r_{\texttt{MB}}=1.000$ vs. $r_{\texttt{HMC}}=0.996$), random effects ($r_{\texttt{MB}}=0.987$ vs. $r_{\texttt{HMC}}=0.974$), and for variance parameters ($r_{\texttt{MB}}=0.997$ vs. $r_{\texttt{HMC}}=0.996$). The same pattern arises for recovery error for fixed effects ($\mathrm{RMSE}_{\texttt{MB}}=0.020$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.197$), random effects ($\mathrm{RMSE}_{\texttt{MB}}=0.097$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.136$), and for variance parameters ($\mathrm{RMSE}_{\texttt{MB}}=0.027$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.029$). Posterior coverage is good for metabeta ($\mathrm{CE}_{\texttt{MB}}=0.021$), whereas HMC’s marginal posterior for the variance parameters tend to be slightly too wide ($\mathrm{CE}_{\texttt{HMC}}=0.146$). Example kernel density estimates of the posteriors for a single regression dataset are plotted in Figure 1C: While the modes often agree, the posterior shapes vary between our model and HMC. Overall, both models perform well on the toy problem with a slight advantage for metabeta. This shows that the pipeline is in principle correctly specified for both approaches.
3.2 In-Distribution Tests
We ran a sweep of tests to check how well our model can handle different mixed-effect datasets in terms of the problem size (determined by $d$ and $q$), numbers of total observations ($n=\sum_{i=1}^{m}n_{i}$) and signal to noise ratio, $\mathrm{SNR}=\mathbb{V}(\mathbf{y}-\bm{\varepsilon})/\mathbb{V}(\bm{\varepsilon})$. The test sets were constructed using the same simulation pipeline as the training sets, but with different random seeds. Table 1 shows that our model’s performance is most strongly affected by the problem size in terms of $\mathrm{RMSE}$. Both relatively low signal to noise ratio and low $n$ affect metabeta’s performance slightly, but none is systematically more impactful than the other. However, its Pearson correlations and coverage are relatively robust to both problem size and dataset properties. HMC, on the other hand, tends to produce more outliers for larger $n$ and $\mathrm{SNR}$. Accordingly, its coverage varies more strongly over datasets. Overall, our model appears to have comparably stable performance and outperforms HMC in most test cases.
@@ 3.2 In-Distribution Tests
metabeta HMC
$d$ $q$ property split $r$ $\mathrm{RMSE}$ $\mathrm{CE}$ $r$ $\mathrm{RMSE}$ $\mathrm{CE}$
@@
$\circ$ 0.995 0.546 -0.071 0.991 1.475 0.012
[…]
$\circ$ 0.991 0.689 -0.021 0.998 0.679 0.129
[…]
$\circ$ 0.968 1.064 -0.085 0.840 5.458 -0.077
[…]
$\circ$ 0.988 0.729 0.009 0.946 1.899 0.090
[…]
$\circ$ 0.975 1.212 -0.078 0.899 2.962 -0.053
[…]
$\circ$ 0.966 1.309 0.020 0.956 1.516 0.086
[…]
$\circ$ 0.961 1.699 -0.082 0.946 1.662 -0.025
[…]
$\circ$ 0.970 1.587 0.014 0.928 3.993 0.067
Table 2: Performance evaluation for metabeta and HMC on semi-synthetic test sets using real predictors $\mathbf{x_{ij}}$ and simulated regression parameters. The evaluation metrics are Pearson’s correlation-coefficient $r$, root mean squared error $\mathrm{RMSE}$, and coverage error $\mathrm{CE}(\alpha)$ averaged over $\alpha\in\{0.05,0.1,0.2,0.32,0.5\}$, separately per type of parameter. Bold formatting indicates better performance.
metabeta HMC
Source Parameters $r$ $\mathrm{RMSE}$ $\mathrm{CE}$ $r$ $\mathrm{RMSE}$ $\mathrm{CE}$
MathAchieve $\bm{\beta}$ 0.999 0.754 0.023 0.992 1.966 0.030
$\bm{\sigma}$ 0.996 0.316 0.010 0.980 0.778 0.164
$\bm{\alpha}$ 0.987 0.832 0.041 0.956 1.518 0.060
Exam $\bm{\beta}$ 0.999 0.554 0.057 0.995 1.548 0.054
$\bm{\sigma}$ 0.994 0.407 0.009 0.996 0.219 0.183
$\bm{\alpha}$ 0.986 0.786 0.031 0.972 1.108 0.099
Gcsemv $\bm{\beta}$ 0.999 0.596 0.007 0.881 6.634 -0.020
$\bm{\sigma}$ 0.988 0.639 -0.009 0.477 30.083 0.071
$\bm{\alpha}$ 0.964 1.430 0.016 0.889 2.434 -0.009
SleepStudy $\bm{\beta}$ 1.000 0.092 0.147 0.996 1.292 0.194
$\bm{\sigma}$ 0.996 0.310 0.050 0.687 9.709 0.169
$\bm{\alpha}$ 0.975 1.065 0.024 0.972 1.166 0.115
3.3 Out-of-Distribution Tests
We gathered 4 canonical datasets that are often used for demonstration purposes of mixed-effects regression: (1) MathAchieve ($d=5,q=1$), available in nlme (Pinheiro et al., 1999), (2) Exam ($d=4,q=1$) and (3) Gcsemv ($d=3,q=1$), both available in mlmRev (Bates & Bolker, 2020), and (4) sleepstudy ($d=2$, $q=2$), available in lme4 (Bates et al., 2015). All contain numerical and categorial predictors, have varying numbers of groups and observations per group, as well as different ranges for $\mathbf{X}$ and $\mathbf{y}$. For each, we built a semi-synthetic test set using real predictors $\mathbf{X}$, simulated regression parameters and resulting outcomes $\mathbf{y}$. Table 2 lists model performance for all datasets for each type of parameter. Our model has the clear advantage: In $11/12$ cases our model has the better correlation with the true parameters, in $12/12$ cases it has the lower $\mathrm{RMSE}$ and in $10/12$ cases it has the better coverage. HMC particularly struggles with the variance parameters, which do not pose a substantial problem for metabeta. Our model again has considerably fewer outliers (Figure 2A) and better coverage (Figure 2B), which is reflected in its credible intervals (Figure 3A). Both models produce appropriate posterior predictive distributions (Figure 3B). Overall, this suggests that the simulated predictors $\mathbf{X}$, on which our model is trained, approximate the structure of realistic datasets well. This makes our model useful for analyzing the type of datasets practitioners care about.
Figure 2: Results based on MathAchieve. Remaining results are depicted in Appendix F. (A) Parameter Recovery. Our model outperforms HMC on average in terms of r, bias and RMSE for all parameter types, and has fewer outliers. (B) Coverage. Our model’s posterior credible intervals are on average more faithfully tuned.
Figure 3: Results based on MathAchieve (A) Credible Intervals. $95\%$ and $50\%$ credible intervals for metabeta and HMC, compared over different parameter values. Note that discrepancies in width are mirrored in the coverage plot of Figure 2B: HMC has poorer coverage for $\beta_{0}$ and $\sigma_{0}$ and its credible intervals are on average wider than metabeta’s for both parameters. The plot for $\beta_{4}$ is omitted due to space constraints. (B) Posterior Predictive. Observed regression outputs (black) plotted against samples from the posterior predictive (colored) and its mean (grey) for both models. Curves based on kernel density estimates over data points, separately for two randomly chosen datasets.

### v2

@@ 2.4 Learning
To calculate the loss for the global parameters, we use the forward Kullback-Leibler divergence between the unknown true posterior $p(\boldsymbol{\vartheta}|\mathbf{s})$ and its flow-based approximation $p_{\Pi}(\boldsymbol{\vartheta}|\mathbf{s}):=p_{n}(\mathbf{z}_{n}|\mathbf{s})$,
$\ell_{\Pi}(\boldsymbol{\vartheta},\mathbf{s})$ $\propto-\mathbb{E}_{\boldsymbol{\vartheta},\mathbf{s}}\left[\log p_{\Pi}(\boldsymbol{\vartheta}|\mathbf{s})\right]$
$=-\mathbb{E}_{\boldsymbol{\vartheta},\mathbf{s}}\left[\log p_{0}(T(\boldsymbol{\vartheta}|\mathbf{s}))+\log\begin{vmatrix}\det J_{T}(\boldsymbol{\vartheta}|\mathbf{s})\end{vmatrix}\right$],
where $T$ and thereby the approximation $p_{\Pi}(\boldsymbol{\vartheta}|\mathbf{s})$ depend on the posterior network. Since the summary $\mathbf{s}$ of data $\mathbf{D}$ is itself depending on the summary network, the end-to-end loss can be written as
$\ell_{\Pi,\Sigma}(\boldsymbol{\vartheta},\mathbf{D})\propto-\mathbb{E}_{\boldsymbol{\vartheta},\mathbf{D}}\left[\log p_{\Pi}(\boldsymbol{\vartheta}|f_{\Sigma}(\mathbf{D}))\right],$
where the last term can again be expanded like above. The objective for the local parameters is completely analogous. We average the local losses over groups and add the result to the global loss, which follows a potential factorization of the joint posterior over both types of regression parameters (see Appendix D). The expectation is approximated by averaging over the batch. Model weights are updated using the Schedule-Free AdamW optimizer (Defazio et al., 2024). We train separate models for different numbers of fixed effects and random effects until convergence, which requires between $10^{5}$ and $10^{6}$ training sets in our case.
@@ 2.5 Post-Hoc Refinement
In the idealized limit of infinite network capacity, neural posterior flexibility, infinite simulations, and perfectly converged optimization, our model would not require any further correction as it would converge on the true posterior. However, in practice these conditions are never fully met. Learning $p_{\Pi}$ by minimizing the forward Kullback-Leibler Divergence naturally forces $p_{\Pi}$ to be positive wherever $p$ is positive, making $p_{\Pi}$ mass-covering (Jerfel et al., 2021). Thus, we can use importance sampling to improve posterior estimation (Tokdar and Kass, 2010; Dax et al., 2023) to correct for inaccuracies of the amortized estimator. For each sample $\boldsymbol{\vartheta}_{k}\sim p_{\Pi}(\boldsymbol{\vartheta}|\mathbf{D})$ we assign an importance weight,
$w_{k}=\frac{p(\mathbf{D}|\boldsymbol{\vartheta}_{k})p(\boldsymbol{\vartheta}_{k})}{p_{\Pi}(\boldsymbol{\vartheta}_{k}|\mathbf{D})},$
@@
Conformal Prediction
Uncertainty quantification is a hallmark of Bayesian inference, making the fidelity of the approximate posterior’s credible intervals a critical concern. Posterior samples can be used to calculate empirical quantiles and thus also intervals that contain $c\%$ of the posterior density. Due to the mass-covering property of $p_{\Pi}$, the learned posteriors tend to be too wide – i.e. the true parameter is inside the $c\%$ credible interval in more than $c\%$ of the cases. This can be quantified with the coverage error
$\mathrm{CE(\alpha)}=\frac{1}{B}\sum_{b=1}^{B}\mathbbm{1}\left(\vartheta^{(b)}\in C_{\alpha}^{(b)}\right)-(1-\alpha),$
which should asymptotically approach 0 under perfect coverage. Too liberal coverage is a commonly known issue of normalizing flows (Chen et al., 2025; Dheur and Taieb, 2025). Conformal prediction (Shafer and Vovk, 2008; Angelopoulos and Bates, 2022; Vovk et al., 2022) is a general-purpose method that constructs distribution-free prediction sets $\hat{C}_{\alpha}$ such that $\mathbb{P}(\boldsymbol{\vartheta}\in\hat{C}_{\alpha})\geq 1-\alpha$. To construct $\hat{C}_{\alpha}$, we use a calibration set to calculate the difference between the true $\boldsymbol{\vartheta}$ and the closest border of the proposed $1-\alpha$ credible interval $C_{\alpha}$. The $1-\alpha$ quantile of these differences is then added to the proposed interval borders, widening them if the value is positive and narrowing them otherwise. Importantly, this does not require retraining but efficiently refines credible intervals post-hoc.
@@ 3 Results
We test our model (1) on toy data with highly constrained parameters and uncorrelated normal data (Section 3.1), (2) in-distribution data with predictors $\mathbf{X}$ sampled from real datasets (Section 3.2), and (3) out-of-distribution data containing subsets of real data where the parameters $\boldsymbol{\vartheta}$ are unknown and the outcomes $\mathbf{y}$ are kept original (Section 3.3). Each test set has a batch size of $128$ regression datasets with varying numbers of observations and signal-to-noise ratios.
[…]
@@ 3.1 Toy Example
To gauge if the pipeline works for both our model and HMC, we first test both on a toy example with $d=2$ and $q=1$, where the observed single predictor is sampled from a standard normal distribution. Result figures can be found in Appendix F. All models reach almost perfect parameter recovery correlation for fixed effects ($r=0.999$ each), variance parameters ($r_{\texttt{MB}}=0.995$ vs. $r_{\texttt{HMC}}=0.991$ vs. $r_{\texttt{VI}}=0.991$), and random effects ($r=0.959$ each). The same pattern arises for recovery error for fixed effects ($\mathrm{RMSE}_{\texttt{MB}}=0.023$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.021$ vs. $\mathrm{RMSE}_{\texttt{VI}}=0.030$), variance parameters ($\mathrm{RMSE}_{\texttt{MB}}=0.022$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.025$ vs. $\mathrm{RMSE}_{\texttt{VI}}=0.034$), and random effects ($\mathrm{RMSE}_{\texttt{MB}}=0.108$ vs. $\mathrm{RMSE}_{\texttt{HMC}}=0.108$ vs. $\mathrm{RMSE}_{\texttt{VI}}=0.109$), Posterior coverage is good for metabeta ($\mathrm{CE}_{\texttt{MB}}=0.007$), whereas HMC’s marginal posterior for the variance parameters tend to be slightly too wide ($\mathrm{CE}_{\texttt{HMC}}=0.094$) and the ones of VI too narrow ($\mathrm{CE}_{\texttt{VI}}=-0.053$). The median posterior NLLs are in the same neighborhood ($\mathrm{NLL}_{\texttt{MB}}=805.1$ vs. $\mathrm{NLL}_{\texttt{HMC}}=829.0$ vs. $\mathrm{NLL}_{\texttt{VI}}=802.6$) and posterior fits are highly correlated ($\mathrm{r}_{\texttt{MB},\texttt{HMC}}=0.940$, $\mathrm{r}_{\texttt{MB},\texttt{VI}}=0.940$, $\mathrm{r}_{\texttt{HMC},\texttt{VI}}=0.999$). Overall, all models perform excellently on the toy problem and make very similar predictions. This shows that the pipeline is in principle correctly specified for each approach.
3.2 Real Predictors, Simulated Parameters
To get a better estimate for model performance under more realistic conditions, we sample predictors ($\mathbf{X}$) from a large set of real datasets, and combine them with synthetically sampled regression parameters ($\boldsymbol{\vartheta}$) to produce regression outcomes ($\mathbf{y}$). The test sets were constructed using the same simulation pipeline as the training sets, but with different random seeds. Please see Appendix A for details. This approach of testing on semi-synthetic datsets has the following considerable benefits over evaluation on purely real data: (1) The regression models are always correctly specified, (2) we know the ground truth parameters and can thus evaluate parameter recovery and coverage, (3) we can compare the results to in-distribution test data and gauge how well the model transfers to realistic predictors (Lueckmann et al., 2021; Ward et al., 2022).
@@ 3.2 Real Predictors, Simulated Parameters
$d$ & $q$ $\mathrm{model}$ $\mathrm{NLL}$ $r$ $\mathrm{RMSE}$ $\mathrm{CE}$ $\mathrm{time}$
3311 MB 456.1 0.987 0.059 0.028 0.01
HMC 423.7 0.987 0.058 0.046 12.48
VI 528.7 0.983 0.089 -0.161 4.49
5522 MB 355.5 0.966 0.079 0.014 0.01
HMC 351.7 0.976 0.067 0.037 13.68
VI 479.8 0.967 0.092 -0.224 9.41
8833 MB 438.3 0.977 0.048 -0.037 0.01
HMC 417.8 0.964 0.092 -0.138 15.59
VI 642.2 0.883 0.405 -0.347 12.75
121244 MB 534.1 0.938 0.106 0.040 0.01
HMC 504.7 0.945 0.099 -0.205 36.96
VI 757.2 0.849 0.511 -0.398 21.58
Figure 2: Model performance for $d=5$ and $q=2$. Results for other regression problems are depicted in Appendix F. (A) Parameter Recovery. Our model reaches similar performance to HMC in terms of $r$, bias and RMSE for all parameter types. (B) Coverage. Our model’s posterior credible intervals are on average more faithfully tuned, whereas the HMC posteriors tend to be unnecessarily broad.
Table 1 shows model performance for hierarchical regression problems with increasing numbers of fixed and random parameters. Recovery and coverage per parameter type are visualized in Figure 2 for $d=5,q=2$. Median model fits of metabeta and HMC are very similar: While HMC generally explains the outcomes better (mean $\Delta\mathrm{NLL}=21.5$), our model outperforms VI in this regard (mean $\Delta\mathrm{NLL}=-156.0$). Over the test batch, the average model fits are highly correlated between metabeta and HMC ($r=0.94$), indicating overall high agreement between both methods. Similarly, parameter recovery is tied with HMC (mean $\Delta\mathrm{r}=-0.001$, mean $\Delta\mathrm{RMSE}=-0.006$). VI performs similarly well for the two smaller problems but struggles more with the latter two, indicating overall unfavorable recovery in comparison to our model, (mean $\Delta\mathrm{r}=0.047$, mean $\Delta\mathrm{RMSE}=-0.201$). Posterior coverage is generally best for metabeta, as its CE is closest to 0 in all cases. In comparison, HMC has unstable coverage ($\mathrm{CE}_{\mathrm{HMC}}\in[-0.205,0.046]$) and VI tends towards too narrow posteriors (mean $\mathrm{CE}_{\mathrm{VI}}=-0.283$). Lastly, the median run time for a single dataset is almost instantaneous for metabeta ($\approx 0.01$s), strongly outperforming both other methods (up to $37$s). Overall, our model appears to have comparably stable performance to HMC and outperforms VI, which marks metabeta as a strong alternative to HMC if practitioners are willing to accept a minor reduction in accuracy for a substantial boost in speed.
3.3 Real Datasets
We gathered 7 canonical datasets that are often used for demonstration purposes of mixed-effects regression, and ran each model on multiple subsets thereof. No parameters are simulated for these test sets and we use the default automatic prior specification of Bambi for posterior estimation. To gauge model fits, we compare in-sample posterior accuracy with the same methods as above. Table 2 lists model performance for all datasets. Overall, the median NLL of our model is very similar to that of HMC and VI, which also shows in the average correlation over batches, ($\mathrm{r}_{\texttt{MB},\texttt{HMC}}=0.880$, $\mathrm{r}_{\texttt{MB},\texttt{VI}}=0.876$, $\mathrm{r}_{\texttt{HMC},\texttt{VI}}=0.951$). This indicates general agreement posterior inference, even on out-of-distribution data with likely miss-specified priors and model structure.
@@ 3.3 Real Datasets
MB HMC VI
Sleep 109.4 115.2 105.3
Gcsemv 390.9 389.2 403.3
Exam 784.8 773.5 782.2
Math 883.2 856.4 869.1
Titanic 810.8 788.7 787.6
Schooling 738.2 640.9 661.9
News 540.3 400.0 399.3